Find the values of the indicated currents and voltage

30 Ω και , - 15 V( + 1, 20Ω 1/4 A Yξ 20Ω Ι. ξ 30Ω 30 Ω 15 & 20Ω -0.5ΑΙ
Answer #1

30 100D 15V 1 0 u - *=c 2ch 15-: - 30 26 30 -20 = 30 5 - 3D (* - by | glis gla → 7+2= TE Res: 30 ×2C = to = · 2014 12人死 | y